OpenStudy (anonymous):

When a question asks me to 'evaluate' does that mean I need to solve the equation ?

hartnn (hartnn):

evaluate means finding the value of expression solve means finding the value of variable in that equation....

OpenStudy (anonymous):

haha I thought you left me !

OpenStudy (anonymous):

Evaluate x2 + y2 when x = 3 and y = 4

OpenStudy (anonymous):

does that mean substitute the x and y with the numbers and add it all up ?

hartnn (hartnn):

yes, correct

hartnn (hartnn):

=3^2+4^2

hartnn (hartnn):

which equals ?

OpenStudy (anonymous):

hang on ! ahaha let me get this down too :p

OpenStudy (anonymous):

25 ?

OpenStudy (anonymous):

or do I need to sqare root the 25 too ?

hartnn (hartnn):

no, its 25 only 3^2+4^2 = 25

OpenStudy (anonymous):

cool :D

hartnn (hartnn):

:)

OpenStudy (anonymous):

Expand (x + 1)(x – 2)

hartnn (hartnn):

( x + 1 ) ( x – 2 )

hartnn (hartnn):

just kidding, u know how to foil ??

OpenStudy (anonymous):

Expanding, no hope. Simplifying, I feel like I have some hope aha

hartnn (hartnn):

ham then try to simplify

OpenStudy (anonymous):

foil ? aha

hartnn (hartnn):